Project

General

Profile

Test Beam Shift Bulletin Board » History » Version 40

Version 39 (Alexandre Pereira Sousa, 07/04/2019 04:02 PM) → Version 40/97 (Alexandre Pereira Sousa, 07/04/2019 04:05 PM)

h1. Test Beam Shift Bulletin Board

Welcome to the NOvA Test Beam shift bulletin board! Look here for relevant information about the current status of operations.

h2. Current Run Coordinator: Alex Sousa (617-682 6064)

h2. Current On-Call Experts:

* Day: Teresa Lackey (608-443-9109)
* Night: Andrew Sutton (919-619-8866)

h2. Current Running Conditions

* Beam on week
* Beamline DAQ on for expert testing, TOF trigger testing, and halo studies
* Run detector DAQ 24/7 depending on halo studies. *If MC7b dew point reaches 9.9 °C, stop detector run, configure APDs warm, and turn off detector HV, then LV, following* [[Expert Procedures]]. Bring detector back up when dew point drops to 8 °C or less and oscillates around that value.
* Information on run plan for this week is found at https://docs.google.com/document/d/1OcYGNOrazkMGkjaUOfP_gn5z7FWEKCR49wYXIb3DNyo/edit?usp=sharing
* *Information for each run must be added by shifter to* https://docs.google.com/spreadsheets/d/1kUI1HRDhCQWzimxRcGTcgThaRhS2VXPaaW3TXKHmLLo/edit?usp=sharing

h2. Secondary Beamline Tuning and Beam Halo Studies

ECL entries with details entry 2497 contains a description of the studies done:
* Day 1: 2497 -
http://dbweb5.fnal.gov:8080/ECL/novatestbeam/E/show?e=2497
*Day 2: 2567 - http://dbweb5.fnal.gov:8080/ECL/novatestbeam/E/show?e=2567

All plots going into: http://novadaq-test-master.fnal.gov/beamtuning/
Running conditions and run numbers found in: https://docs.google.com/spreadsheets/d/1kUI1HRDhCQWzimxRcGTcgThaRhS2VXPaaW3TXKHmLLo/edit?usp=sharing
Excellent Evernote summary of studies by Dung Phan:
* Day 1: https://www.evernote.com/l/AOdmsFKXTUxG2qptCTOJQ1AV3UtCIpyhc6k
* Day2: https://www.evernote.com/l/AOezRbwJmtFP6JdVqmmlgHK2rMhYO3MNFQU

h2. Current DAQ Configurations

h3. Beamline

Two conditions:

on: TOFPROMPT MWPCOFF SWYDGATE
off: DIGBUSY

on: TOFUS MWPCOFF SWYDGATE
off: DIGBUSY

h3. Detector

* Configuration: TestBeam_AllTrigs_HalfDetector

h2. Current Digitizer Configurations

* Digitizer rate: 2.5 GHz

h2. Important Notes

* Every time a beamline run ends, or at the end of your shift, please add it to the table at [[Commissioning runs]].

* The detector is now interlocked with environmental monitoring, so if there is an environment alarm, the HV and LV power supplies will shutdown, preventing damage to APDs and FEBs. However, you should keep an eye on the "PLC Dry Gas & Water" synoptic display, in case the HVAC freezes or the water chiller trips. If the HVAC stops cooling MC7 (most likely triggering an environment dew point alarm), the water chiller should be stopped, to prevent condensation on the water lines flowing through the detector, until the HVAC is recovered. If the water chiller stops working (most likely triggering an Inlet APD Water Temp alarm), a picture of the front panel in the water chiller should be taken, and the water chiller should be stopped until it can be recovered. More information is found under: [[Environment Emergency Procedures]]

h2. Temporary Things To Be Aware Of!

* If the VNC crashes, it can be reopened from the icon on the desktop. This will pop up a terminal and open up the display. If the terminal doesn't close on its own, please close it yourself otherwise the VNC may not start properly.

* We are still confirming the beamline TDU is working as expected. Things may not be too smooth with this.

* The beamline DQM is still a little flaky, and we're working on it. Try restarting a few times if it crashes, it normally works. If you can't get it to work, and you're concerned about the quality of the data, contact the on-call expert.